# An Algebraic Roadmap of Particle Theories Part I: General construction

### v1

@@ Abstract
It is a curious fact that our most well-studied BSM theories can inform us about the Standard Model itself. Just over a decade ago, [1], Baez and Huerta explained how the Standard Model’s gauge group can be seen to coincide with the intersection of Georgi and Glashow’s SU(5), and Pati and Salam’s $SU(4)\times SU(2)\times SU(2)$, inside Spin(10). More recently, [2], [3], it was shown how (up to a factor of B-L) the Spin(10) model, the Pati-Salam model, the Left-Right Symmetric model, the Standard Model pre-Higgs mechanism, and the Standard Model post-Higgs mechanism are interlinked via a sequence of division algebraic reflections. Between these three papers, [1], [2], [3], are six theories that have been studied extensively by particle physicists. In this article, we demonstrate how the full set of six familiar particle models may be laid out in one connected algebraic particle roadmap.
The inclusion of a quaternionic reflection within the network further differentiates $W^{\pm}$ bosons from the $Z^{0}$ boson in comparison to the Standard Model. It may introduce subtle new considerations for the phenomenology of electroweak symmetry breaking.
@@ I Introduction
Since at least as early as the 1970s, physicists and mathematicians have been identifying Clifford algebraic and division algebraic patterns within the architecture of elementary particles. Now half a century ago, Günaydin and Gürsey proposed a quark model based on octonions, [6]. Shortly thereafter came a series of papers by Casalbuoni et al who proposed particle models based on a large variety of Clifford algebras, [7], [8], [9]. Since these days, many authors have invested a great deal of time and effort to the endeavour. In what is apologetically far from an exhaustive list, we point out the Clifford algebraic work of Trayling and Baylis (Cl(0,7)) [10], Barrett (via NCG) [11], Zenczykowski ($\mathbb{C}l(6)$) [12], Connes et al (via NCG) [13], Gording and Schmidt-May ($\mathbb{C}l(6)$) [14], Todorov ($\mathbb{C}l(10)$) [15], Borštnik (Cl(p,q)) [16], the division algebraic work Adler ($\mathbb{H}$) [17], De Leo ($\mathbb{H}$) [18], Toppan et al ($\mathbb{H},\mathbb{O}$) [19], Baez ($\mathbb{R},\mathbb{C},\mathbb{H}$) [20], Duff et al ($\mathbb{R},\mathbb{C},\mathbb{H},\mathbb{O}$) [21], Hughes ($\mathbb{O}$) [22], Catto et al ($\mathbb{O}$) [23], Gresnigt ($\mathbb{O}$) [24], Asselmeyer-Maluga ($\mathbb{H},\mathbb{O}$) [25], Bolokhov ($\mathbb{H}$) [26], Vaibhav and Singh ($-\mathbb{H}$ and split-$\mathbb{O}$) [27], Boyle ($J_{3}(\mathbb{O})$) [28], Jackson ($\mathbb{O}$) [29], Hunt ($\mathbb{R},\mathbb{C},\mathbb{H},\mathbb{O}$) [30], Lasenby ($\mathbb{O}$) [31], Manogue et al ($E_{8}$) [32], [33], Hun Jang, (Hypercomplex) [35], Hiley ($-\mathbb{H}$) [36], and most recently Penrose (split-$\mathbb{O}$) [34], the $\mathbb{R}\otimes\mathbb{C}\otimes\mathbb{H}\otimes\mathbb{O}$ work of Dixon, [37], Perelman, [38], Chester et al, [39], Köplinger, [40].
@@ II.1 Algebra of grand unified theories
$\begin{array}[]{ccc}\mathfrak{so}(10)&\supset&\mathfrak{su}(5)\\ \mathbf{16}&\hskip 28.45274pt\mapsto&\mathbf{10}\oplus\mathbf{5^{*}}\oplus\mathbf{1}\end{array}$ (1)
@@
In the same era, another (partially) unified theory was constructed by Pati and Salam, [45]. Their intention was to capitalize on an observed pattern-matching between quarks and leptons. Via the group $SU(4)\times SU(2)\times SU(2)$, the Standard Model’s three “red, green, blue” quarks colours were augmented to include a fourth “lilac” lepton colour. Furthermore, a symmetry between left- and right-handed particles was conjectured. Curiously enough, Pati and Salam’s model also fits neatly into the $\mathbf{16}$ of Spin(10):
$\begin{array}[]{ccc}\mathfrak{so}(10)&\supset&\mathfrak{su}(4)\oplus\mathfrak{su}(2)\oplus\mathfrak{su}(2)\\ \mathbf{16}&\hskip 28.45274pt\mapsto&\left(\mathbf{4},\mathbf{2},\mathbf{1}\right)\oplus\left(\mathbf{4^{*}},\mathbf{1},\mathbf{2}\right)\end{array}$ (2)
@@ II.2 Division algebraic symmetry breaking
$\begin{array}[]{ccc}\mathfrak{so}(10)&\mapsto&\mathfrak{so}(6)\oplus\mathfrak{so}(4)\\ \mathbf{10}&\hskip 28.45274pt\mapsto&\mathbf{6}\oplus\mathbf{4}\end{array}$ (3)
derives from the octonions was first discovered by Furey and Hughes.
These results were distributed widely amongst colleagues on the 16th of February, 2021, [2], [3], and discussed extensively with some of them, [46]. In particular, Kirill Krasnov contributed to our efforts by following suit and reproducing our results, thus confirming them.
This octonionic structure may be seen to be responsible for sending the Spin(10) model to the Pati-Salam model, and the SU(5) model to the Standard Model.
@@ 1. An additional B-L symmetry persisted,
In this article, we demonstrate that both of these issues may be addressed simultaneously. Furthermore, we offer a concise answer to those challenges posed by Baez and Huerta as to where from their mysterious conditions may be seen to originate.
As will be seen, the overarching theme may be described as the implementation of simultaneous group actions on the fermionic representations. These simultaneous group actions, or more precisely, Lie algebra actions, allow us to merge the diagrams of Figures 2 and 3. The result (Figure 6) is a detailed roadmap of particle models.
@@ II.3 A consistency condition
In this article, we entertain the idea that Standard Model fermions should simultaneously transform under both the known $\mathfrak{spin}(2n)$ spinor action, and also under the known $\mathfrak{so}(2n)$ multivector action. This consistency condition reduces $\mathfrak{spin}(2n)$ to $\mathfrak{su}(n)$. In one stroke, this single requirement addresses both the chirality and B-L issues plaguing [3]. It also fuses Baez and Huerta’s conditions 2. and 3. in Section II.1 into a single constraint. It may be seen to be algebraically equivalent to a statement made in a footnote of [9], where a volume form is fixed. Here, we derive the result instead as a consistency condition in the construction of fermions as multivectors.
@@ III Simultaneous group actions of the multivector type
$\begin{array}[]{rcl}\textup{Spin(10)}&\mapsto&\textup{SU(5),}\\ \textup{Pati-Salam}&\mapsto&\textup{Standard Model,}\\ \textup{Left-Right Symmetric}&\mapsto&\textup{Standard Model.}\end{array}$
@@ III.2 Minimal left ideal construction
$\begin{array}[]{l}\{a_{j_{1}},a_{j_{2}}\}=\{a_{j_{1}}^{\dagger},a_{j_{2}}^{\dagger}\}=0\\ \{a_{j_{1}},a^{\dagger}_{j_{2}}\}=\delta_{j_{1}j_{2}}\end{array}$ (7)
@@
$\begin{array}[]{lc}\Psi=&c_{12\dots n}\hskip 1.42262pta_{1}a_{2}\cdots a_{n}v\\ &\vdots\\ &+\hskip 2.84526ptc_{12}\hskip 1.42262pta_{1}a_{2}v+c_{13}\hskip 1.42262pta_{1}a_{3}v\dots+c_{n-1\hskip 1.42262ptn}\hskip 1.42262pta_{n-1}a_{n}v\\ &+\hskip 2.84526ptc_{1}\hskip 1.42262pta_{1}v+c_{2}\hskip 1.42262pta_{2}v\dots+c_{n}\hskip 1.42262pta_{n}v\\ &+\hskip 2.84526ptc_{0}\hskip 1.42262ptv\end{array}$ (10)
for $c_{0},c_{j},\dots,c_{12\dots n}\in\mathbb{C}$. From here, it is straightforward to to see that $\Psi$ takes the form of a Fock space, familiar to physicists.
@@ III.4 $\mathfrak{spin}(2n)$ $\mapsto$ $\mathfrak{su}(n)$ via the
Equation (14) supplies a derivation. One would expect an induced infinitesimal transformation of $\Psi$ to go as
@@
$\begin{array}[]{c}\delta\Psi=[r_{k_{1}k_{2}}\Gamma_{k_{1}}\Gamma_{k_{2}},\hskip 1.42262pt\Psi]\hskip 2.84526pt=r_{k_{1}k_{2}}\Gamma_{k_{1}}\Gamma_{k_{2}}\Psi\\ \Rightarrow\\ \mathfrak{so}(2n)\hskip 5.69054pt\mapsto\hskip 5.69054pt\mathfrak{su}(n).\end{array}$ (18)
@@
$\begin{array}[]{lll}\ell_{n}&=&R_{j_{1}j_{2}}\left(a_{j_{1}}a^{\dagger}_{j_{2}}-a_{j_{2}}a^{\dagger}_{j_{1}}\right)\\ &+&R_{j_{1}j_{2}}^{\prime}i\left(a_{j_{1}}a^{\dagger}_{j_{2}}+a_{j_{2}}a^{\dagger}_{j_{1}}\right)\\ &+&R_{j}\hskip 1.42262pti\left(a_{j}a^{\dagger}_{j}-a_{j+1}a^{\dagger}_{j+1}\right),\end{array}$ (19)
@@
$\begin{array}[]{lll}\ell_{n}&=&r_{j}\hskip 1.42262pt\Gamma_{j}\Gamma_{j+n}\\ &+&r_{j_{1}j_{2}}\left(\Gamma_{j_{1}}\Gamma_{j_{2}}+\Gamma_{j_{1}+n}\Gamma_{j_{2}+n}\right)\\ &+&r_{j_{1}j_{2}}^{\prime}\left(\Gamma_{j_{1}}\Gamma_{j_{2}+n}+\Gamma_{j_{2}}\Gamma_{j_{1}+n}\right),\end{array}$ (20)
@@ III.5 Significance
$\begin{array}[]{rcl}\textup{Spin(10)}&\mapsto&\textup{SU(5),}\\ \textup{Pati-Salam}&\mapsto&\textup{Standard Model,}\\ \textup{Left-Right Symmetric}&\mapsto&\textup{Standard Model,}\end{array}$
this consistency condition supplies a natural explanation for the Standard Model’s infamous chirality.
@@ IV.1 Motivation for considering $\mathbb{R}\otimes\mathbb{C}\otimes\mathbb{H}\otimes\mathbb{O}$
$\begin{array}[]{l}\textup{Spin(10)}\mapsto\textup{Pati-Salam}\mapsto\textup{Left-Right Symmetric}\mapsto\\ \hskip 21.33955pt\textup{Pre-Higgs Standard Model (+ B-L)}\mapsto\\ \hskip 42.67912pt\textup{Post-Higgs Standard Model (+ B-L).}\end{array}$
@@ IV.3 From $\mathbb{R}\otimes\mathbb{C}\otimes\mathbb{H}\otimes\mathbb{O}$ to its space of linear operators
First let us define what we mean by multiplication algebra. Suppose $x,y$ are elements in an algebra $\mathbb{D}.$ We then define $L_{x}(y):=xy$. Similarly, define $R_{x}(y):=yx$. Both $L_{x}$ and $R_{x}$ may then be seen as (possibly distinct) linear maps sending $y$ to some new element in $\mathbb{D}.$ Hence, $L_{x}$ and $R_{x}$ $\in End(\mathbb{D}).$
@@ IV.4 Clifford factors
$\begin{array}[]{lllll}&&\mathcal{L}_{\mathbb{C}}=\mathcal{R}_{\mathbb{C}}&\hskip 5.69054pt\leftrightarrow&Cl\left(0,1\right)\\ &&\mathcal{L}_{\mathbb{H}}&\hskip 5.69054pt\leftrightarrow&Cl\left(0,2\right)\\ &&\mathcal{R}_{\mathbb{H}}&\hskip 5.69054pt\leftrightarrow&Cl\left(0,2\right)\\ &&\mathcal{L}_{\mathbb{O}}\simeq\mathcal{R}_{\mathbb{O}}&\hskip 5.69054pt\leftrightarrow&Cl\left(0,6\right).\\ \end{array}$ (21)
@@
$\begin{array}[]{lll}R_{e_{7}}f:=fe_{7}&=&\frac{1}{2}\left(-e_{7}f+e_{1}\left(e_{3}f\right)+e_{2}\left(e_{6}f\right)+e_{4}\left(e_{5}f\right)\right)\\ &=&\frac{1}{2}\left(-L_{e_{7}}+L_{e_{1}}L_{e_{3}}+L_{e_{2}}L_{e_{6}}+L_{e_{4}}L_{e_{5}}\right)f,\end{array}$ (22)
@@
$\begin{array}[]{l}\mathcal{L}_{\mathbb{C}}\otimes\mathcal{L}_{\mathbb{H}}\otimes\mathcal{L}_{\mathbb{O}}\\ \simeq Cl(0,1)\otimes Cl(0,2)\otimes Cl(0,6)\simeq\mathbb{C}l(8).\end{array}$ (23)
@@
$\begin{array}[]{l}\mathcal{L}_{\mathbb{C}}\otimes\mathcal{L}_{\mathbb{H}}\otimes\mathcal{R}_{\mathbb{H}}\otimes\mathcal{L}_{\mathbb{O}}\\ \simeq Cl(0,1)\otimes Cl(0,2)\otimes Cl(0,2)\otimes Cl(0,6)\simeq\mathbb{C}l(10).\end{array}$ (24)

### v2

@@ Abstract
Expanding the results of [1], [2], [3], we demonstrate a network of algebraic connections between six well-known particle theories. These are the Spin(10) model, the Georgi-Glashow model, the Pati-Salam model, the Left-Right Symmetric model, the Standard Model both pre- and post-Higgs mechanism.
A new inclusion of a quaternionic reflection within the network further differentiates $W^{\pm}$ bosons from the $Z^{0}$ boson in comparison to the Standard Model. It may introduce subtle new considerations for the phenomenology of electroweak symmetry breaking.
@@ I Introduction
Since at least as early as the 1970s, physicists and mathematicians have been identifying Clifford algebraic and division algebraic patterns within the architecture of elementary particles. Now half a century ago, Günaydin and Gürsey proposed a quark model based on octonions, [6]. Shortly thereafter came a series of papers by Casalbuoni et al who proposed particle models based on a large variety of Clifford algebras, [7], [8], [9]. Since these days, many authors have invested a great deal of time and effort to the endeavour. In what is apologetically far from an exhaustive list, we point out the Clifford algebraic work of Trayling and Baylis (Cl(0,7)) [10], Barrett (via NCG) [11], Zenczykowski ($\mathbb{C}l(6)$) [12], Connes et al (via NCG) [13], Stoica ($\mathbb{C}l(6)$) [14], Gording and Schmidt-May ($\mathbb{C}l(6)$) [15], Todorov ($\mathbb{C}l(10)$) [16], [17], Borštnik (Cl(p,q)) [18], the division algebraic work of Conway ($\mathbb{C}\otimes\mathbb{H}$) [19], Silagadze ($J_{3}(\mathbb{O})$) [20], Adler ($\mathbb{H}$) [21], De Leo ($\mathbb{H}$) [22], Toppan et al ($\mathbb{H},\mathbb{O}$) [23], Baez ($\mathbb{R},\mathbb{C},\mathbb{H}$) [24], Duff et al ($\mathbb{R},\mathbb{C},\mathbb{H},\mathbb{O}$) [25], Hughes ($\mathbb{O}$) [26], Catto et al ($\mathbb{O}$) [27], Gresnigt ($\mathbb{O}$) [28], Asselmeyer-Maluga ($\mathbb{H},\mathbb{O}$) [29], Bolokhov ($\mathbb{H}$) [30], Vaibhav and Singh ($-\mathbb{H}$ and split-$\mathbb{O}$) [31], Boyle ($J_{3}(\mathbb{O})$) [32], Jackson ($\mathbb{O}$) [33], Hunt ($\mathbb{R},\mathbb{C},\mathbb{H},\mathbb{O}$) [34], Lasenby ($\mathbb{O}$) [35], Manogue et al ($E_{8}$) [36], [37], Hun Jang, (Hypercomplex) [39], Hiley ($-\mathbb{H}$) [40], and most recently Penrose (split-$\mathbb{O}$) [38], the $\mathbb{R}\otimes\mathbb{C}\otimes\mathbb{H}\otimes\mathbb{O}$ work of Dixon, [41], [42], Castro Perelman, [43], Chester et al, [44], Köplinger, [45].
In earlier years, several authors have employed the division algebras in order to break a variety of symmetries. In [6], Günaydin and Gürsey made use of an octonionic imaginary unit so as to break $G_{2}\mapsto$ SU(3) in the context of a quark model. In [41], Dixon made use of two octonionic projection operators in order to reduce Spin(1,9)$\times$×SU(2) to Spin(1,3) and a non-chiral representation of the Standard Model gauge group. These groups acted on a fermionic space described by two copies of $\mathbb{R}\otimes\mathbb{C}\otimes\mathbb{H}\otimes\mathbb{O},$ later to be known as the Dixon algebra. In [32], Boyle studied an $E_{6}$ model in the context of a complexified version of the exceptional Jordan algebra, $J_{3}(\mathbb{O})$. There, he found that a single octonionic imaginary unit may break Spin(10)$\subset E_{6}$ to the Left-Right symmetric model.
Other related symmetry breaking steps were described in the 1970s by Casalbuoni and Gatto in a footnote of [9]. There a Spin(2n) $\mapsto$ SU(n) breaking is effected by requiring the invariance of a fermionic monomial. Similarly, in [10], Trayling and Baylis proposed fixing the sterile neutrino in the context of Cl(0,7). Subsequently Todorov also proposed fixing the sterile neutrino in [16] in the context of $\mathbb{C}l(10)$. These symmetry breaking steps are closely tied to Baez and Huerta’s proposal to fix a fermionic volume element in [1].
@@ II.1 Algebra of grand unified theories
$\begin{array}[]{ccc}\mathfrak{so}(10)&\supset&\mathfrak{su}(5)\vspace{2mm}\\ \mathbf{16}&\hskip 28.45274pt\mapsto&\mathbf{10}\oplus\mathbf{5^{*}}\oplus\mathbf{1}\end{array}$ (1)
@@
In the same era, another (partially) unified theory was constructed by Pati and Salam, [50]. Their intention was to capitalize on an observed pattern-matching between quarks and leptons. Via the group SU(4)$\times$×SU(2)$\times$×SU(2), the Standard Model’s three “red, green, blue” quarks colours were augmented to include a fourth “lilac” lepton colour. Furthermore, a symmetry between left- and right-handed particles was conjectured. Curiously enough, Pati and Salam’s model also fits neatly into the $\mathbf{16}$ of Spin(10):
$\begin{array}[]{ccc}\mathfrak{so}(10)&\supset&\mathfrak{su}(4)\oplus\mathfrak{su}(2)\oplus\mathfrak{su}(2)\vspace{2mm}\\ \mathbf{16}&\hskip 28.45274pt\mapsto&\left(\mathbf{4},\mathbf{2},\mathbf{1}\right)\oplus\left(\mathbf{4^{*}},\mathbf{1},\mathbf{2}\right)\end{array}$ (2)
@@ II.2 Division algebraic symmetry breaking
$\begin{array}[]{ccc}\mathfrak{so}(10)&\mapsto&\mathfrak{so}(6)\oplus\mathfrak{so}(4)\vspace{2mm}\\ \mathbf{10}&\hskip 28.45274pt\mapsto&\mathbf{6}\oplus\mathbf{4}\end{array}$ (3)
results from preserving an octonionic volume element was first discovered in [3]. This octonionic structure may be seen to be responsible for sending the Spin(10) model to the Pati-Salam model, and the SU(5) model to the Standard Model.
@@ 1. An additional B-L symmetry persisted,
In this article, both of these issues are addressed simultaneously. Furthermore, a concise answer s offered to those challenges posed by Baez and Huerta as to where from their mysterious conditions may be seen to originate.
As will be seen, the overarching theme may be described as the implementation of simultaneous group actions on the fermionic representations. These simultaneous group actions, or more precisely, Lie algebra actions, allow one to upgrade and merge the diagrams of Figures 2 and 3. The result (Figure 6) is a detailed roadmap of particle models.
@@ II.3 A consistency condition
In this article, we entertain the idea that Standard Model fermions should simultaneously transform under both the known $\mathfrak{spin}(2n)$ spinor action, and also under the known $\mathfrak{so}(2n)$ multivector action. This consistency condition reduces $\mathfrak{spin}(2n)$ to $\mathfrak{su}(n)$. In one stroke, this single requirement addresses both the chirality and B-L issues plaguing [3]. It also fuses Baez and Huerta’s conditions 2. and 3. in Section II.1 into a single constraint. Here, we derive the result as a consistency condition in the construction of fermions as multivectors.
@@ III Simultaneous group actions of the multivector type
$\begin{array}[]{rcl}\textup{Spin(10)}&\mapsto&\textup{SU(5),}\vspace{2mm}\\ \textup{Pati-Salam}&\mapsto&\textup{Standard Model,}\vspace{2mm}\\ \textup{Left-Right Symmetric}&\mapsto&\textup{Standard Model.}\end{array}$
@@ III.2 Minimal left ideal construction
$\begin{array}[]{l}\{a_{j_{1}},a_{j_{2}}\}=\{a_{j_{1}}^{\dagger},a_{j_{2}}^{\dagger}\}=0\vspace{2mm}\\ \{a_{j_{1}},a^{\dagger}_{j_{2}}\}=\delta_{j_{1}j_{2}}\end{array}$ (7)
@@
$\begin{array}[]{lc}\Psi=&z_{12\dots n}\hskip 1.42262pta_{1}a_{2}\cdots a_{n}v\vspace{2mm}\\ &\vdots\vspace{2mm}\\ &+\hskip 2.84526ptz_{12}\hskip 1.42262pta_{1}a_{2}v+z_{13}\hskip 1.42262pta_{1}a_{3}v\dots+z_{n-1\hskip 1.42262ptn}\hskip 1.42262pta_{n-1}a_{n}v\vspace{2mm}\\ &+\hskip 2.84526ptz_{1}\hskip 1.42262pta_{1}v+z_{2}\hskip 1.42262pta_{2}v\dots+z_{n}\hskip 1.42262pta_{n}v\vspace{2mm}\\ &+\hskip 2.84526ptz_{0}\hskip 1.42262ptv\end{array}$ (10)
for $z_{0},z_{j},\dots,z_{12\dots n}\in\mathbb{C}$. From here, it is straightforward to see that $\Psi$ takes the form of a Fock space, familiar to physicists.
@@ III.4 $\mathfrak{spin}(2n)$ $\mapsto$ $\mathfrak{su}(n)$ via the
Equation (14) supplies a derivation. One would expect an induced infinitesimal transformation of $\Psi$ to materialize as
@@
$\begin{array}[]{c}\delta\Psi=[r_{k_{1}k_{2}}\Gamma_{k_{1}}\Gamma_{k_{2}},\hskip 1.42262pt\Psi]\hskip 2.84526pt=r_{k_{1}k_{2}}\Gamma_{k_{1}}\Gamma_{k_{2}}\Psi\vspace{2mm}\\ \Rightarrow\vspace{2mm}\\ \mathfrak{so}(2n)\hskip 5.69054pt\mapsto\hskip 5.69054pt\mathfrak{su}(n).\end{array}$ (18)
@@
$\begin{array}[]{lll}\ell_{n}&=&R_{j_{1}j_{2}}\left(a_{j_{1}}a^{\dagger}_{j_{2}}-a_{j_{2}}a^{\dagger}_{j_{1}}\right)\vspace{2mm}\\ &+&R_{j_{1}j_{2}}^{\prime}i\left(a_{j_{1}}a^{\dagger}_{j_{2}}+a_{j_{2}}a^{\dagger}_{j_{1}}\right)\vspace{2mm}\\ &+&R_{j}\hskip 1.42262pti\left(a_{j}a^{\dagger}_{j}-a_{j+1}a^{\dagger}_{j+1}\right),\end{array}$ (19)
@@
$\begin{array}[]{lll}\ell_{n}&=&r_{j}\hskip 1.42262pt\Gamma_{j}\Gamma_{j+n}\vspace{2mm}\\ &+&r_{j_{1}j_{2}}\left(\Gamma_{j_{1}}\Gamma_{j_{2}}+\Gamma_{j_{1}+n}\Gamma_{j_{2}+n}\right)\vspace{2mm}\\ &+&r_{j_{1}j_{2}}^{\prime}\left(\Gamma_{j_{1}}\Gamma_{j_{2}+n}+\Gamma_{j_{2}}\Gamma_{j_{1}+n}\right),\end{array}$ (20)
@@ III.5 Significance
$\begin{array}[]{rcl}\textup{Spin(10)}&\mapsto&\textup{SU(5),}\vspace{2mm}\\ \textup{Pati-Salam}&\mapsto&\textup{Standard Model,}\vspace{2mm}\\ \textup{Left-Right Symmetric}&\mapsto&\textup{Standard Model,}\end{array}$
this consistency condition supplies a natural explanation for the Standard Model’s maximal chirality.
@@ IV.1 Motivation for considering $\mathbb{R}\otimes\mathbb{C}\otimes\mathbb{H}\otimes\mathbb{O}$
$\begin{array}[]{l}\textup{Spin(10)}\mapsto\textup{Pati-Salam}\mapsto\textup{Left-Right Symmetric}\mapsto\vspace{2mm}\\ \hskip 21.33955pt\textup{Pre-Higgs Standard Model (+ B-L)}\mapsto\vspace{2mm}\\ \hskip 42.67912pt\textup{Post-Higgs Standard Model (+ B-L).}\end{array}$
@@ IV.3 From $\mathbb{R}\otimes\mathbb{C}\otimes\mathbb{H}\otimes\mathbb{O}$ to its space of linear operators
First let us define what is meant by multiplication algebra. Suppose $x,y$ are elements in an algebra $\mathbb{D}.$ We then define $L_{x}(y):=xy$. Similarly, define $R_{x}(y):=yx$. Both $L_{x}$ and $R_{x}$ may then be seen as (possibly distinct) linear maps sending $y$ to some new element in $\mathbb{D}.$ Hence, $L_{x}$ and $R_{x}$ $\in End(\mathbb{D}).$
@@ IV.4 Clifford factors
$\begin{array}[]{lllll}&&\mathcal{L}_{\mathbb{C}}=\mathcal{R}_{\mathbb{C}}&\hskip 5.69054pt\leftrightarrow&Cl\left(0,1\right)\vspace{2mm}\\ &&\mathcal{L}_{\mathbb{H}}&\hskip 5.69054pt\leftrightarrow&Cl\left(0,2\right)\vspace{2mm}\\ &&\mathcal{R}_{\mathbb{H}}&\hskip 5.69054pt\leftrightarrow&Cl\left(0,2\right)\vspace{2mm}\\ &&\mathcal{L}_{\mathbb{O}}\simeq\mathcal{R}_{\mathbb{O}}&\hskip 5.69054pt\leftrightarrow&Cl\left(0,6\right).\\ \end{array}$ (21)
@@
$\begin{array}[]{lll}R_{e_{7}}f:=fe_{7}&=&\frac{1}{2}\left(-e_{7}f+e_{1}\left(e_{3}f\right)+e_{2}\left(e_{6}f\right)+e_{4}\left(e_{5}f\right)\right)\vspace{2mm}\\ &=&\frac{1}{2}\left(-L_{e_{7}}+L_{e_{1}}L_{e_{3}}+L_{e_{2}}L_{e_{6}}+L_{e_{4}}L_{e_{5}}\right)f,\vspace{2mm}\end{array}$ (22)
@@
$\begin{array}[]{l}\mathcal{L}_{\mathbb{C}}\otimes\mathcal{L}_{\mathbb{H}}\otimes\mathcal{L}_{\mathbb{O}}\vspace{2mm}\\ \simeq Cl(0,1)\otimes Cl(0,2)\otimes Cl(0,6)\simeq\mathbb{C}l(8).\end{array}$ (23)
@@
$\begin{array}[]{l}\mathcal{L}_{\mathbb{C}}\otimes\mathcal{L}_{\mathbb{H}}\otimes\mathcal{R}_{\mathbb{H}}\otimes\mathcal{L}_{\mathbb{O}}\vspace{2mm}\\ \simeq Cl(0,1)\otimes Cl(0,2)\otimes Cl(0,2)\otimes Cl(0,6)\simeq\mathbb{C}l(10).\vspace{2mm}\end{array}$ (24)